The first 3 seconds the car moves with 6m/s speed other 2 seconds the car moves with 14m/s. What is the average speed of the car?
1
Expert's answer
2012-10-24T09:44:59-0400
The car moved the distance
d = 3s * 6m/s + 2s * 14m/s = 18+28 = 46
m
during
t = 3s + 2s = 5s

Hence, by definition, the average speed
is

v = total_distance / total_time = d/t = 46/5 = 9.2 m/s